# Fine-Tuning Flow Matching via Maximum Likelihood Estimation of Reconstructions

### v1

@@ B.4 Proof of Theorem 4
Proof. It follows from the Lipschitz continuity Assumption 3 that $|p_{j}^{i}(x,\xi)|=|f_{j}^{i}((W_{j}x)^{i}+(W_{j}\xi)^{i})-f_{j}^{i}((W_{j}x)^{i})|\leq L_{j}^{i}|(W_{j}\xi)^{i}|$. Here, the superscript $i$ denotes the $i$-th component of the vector. When $(W_{j}x)^{i}\geq 0$, we have $p_{j}^{i}(x,\xi)=f_{j}^{i}((W_{j}x)^{i}+(W_{j}\xi)^{i})-f_{j}^{i}((W_{j}x)^{i})\geq 0$ due to the Monotonicity in Assumption 3. Then $p_{j}^{i}(x,\xi)\leq L_{j}^{i}(W_{j}\xi)^{i}$. Multiplying both sides by a non-negative number $p_{j}^{i}(x,\xi)$, we get $p_{j}^{i}(x,\xi)^{2}\leq L_{j}^{i}p_{j}^{i}(x,\xi)(W_{j}x)^{i}$. When $(W_{j}x)^{i}\leq 0$, we have $p_{j}^{i}(x,\xi)=f_{j}^{i}((W_{j}x)^{i}+(W_{j}\xi)^{i})-f_{j}^{i}((W_{j}x)^{i})\leq 0$ due to the same Monotonicity Assumption, leading to $p_{j}^{i}(x,\xi)\leq-L_{j}^{i}(W_{j}\xi)^{i}$. Multiplying both sides by a non-negative number $p_{j}^{i}(x,\xi)$, we get the same result $p_{j}^{i}(x,\xi)^{2}\leq L_{j}^{i}p_{j}^{i}(x,\xi)(W_{j}x)^{i}$.
Summing over $i$ on both sides gives $\sum_{i}p_{j}^{i}(x,\xi)^{2}\leq\sum_{i}L_{j}^{i}p_{j}^{i}(x,\xi)(W_{j}x)^{i}$. Or equivalently, in a compact from, $p_{j}(x,\xi)^{\top}p_{j}(x,\xi)\leq(W_{j}x)^{\top}L_{j}p_{j}(x,\xi)=x^{\top}W_{j}^{\top}L_{j}p_{j}(x,\xi)$. This completes the proof of the first part of the lemma. By noting the relation of $f(W_{j}x)=f(W_{j}x-0)=p_{j}(0,\xi)$, the second part of the lemma holds naturally.
Proof (Proof of Theorem 4).
@@
Then, under Lemma 2, it can be deduced that
@@
Therefore, with the conditions equation 10, we can substantiate that the error dynamics of system equation 7 asymptotically approaches zero, meaning that the solution is contracting. Moreover, by Khalil & Grizzle (2002, Theorem 4.9). the stablility area of $\xi$ contains $\{\xi|V(\xi)\leq\max_{\zeta\in\mathbb{R}^{d}}\tilde{V}(\zeta)\}$. Since $y(t)=x(t)+\xi(t)$, we can determine that $\forall y_{0}\in\{x_{0}+\xi|V(\xi)\leq\max_{\zeta\in\mathbb{R}^{d}}\tilde{V}(\zeta)\}$, $y(t)\to x(t)$. This completes the proof.
@@
Appendix C More Discussion on Reconstruction Error Optimization in Precision-Demanding Tasks
@@
Appendix D Experiment Details and Additional Results
D.1 cifar-10
We perform an experiment on unconditional CIFAR-10 generation from a Gaussian source to examine how OT-CFM performs in the high-dimensional image setting. We use a similar setup to that of Lipman et al. (2023) and Tong et al. (2024), including the time-dependent U-Net architecture from Nichol & Dhariwal (2021) that is commonly used in diffusion models. We choose channels =128=128, depth =2=2, channels multiple =[1,2,2,2]=[1,2,2,2], heads =4=4, heads channels =64=64, attention resolution =16=16, dropout =0.1=0.1, batch size per gpu =128=128, gpus =4=4, epochs =2000=2000. We use a constant learning rate, set to $2\times 10^{-4}$ in the pre-training stage and $5\times 10^{-6}$ in the fine-tuning stage. To prevent training instabilities and variance, we clip the gradient norm to 1 and rely on exponential moving average with a decay of 0.99.
The visualization results of the generated images are presented in Fig. E.2. From Figs 2(b)-2(b), we can see that fine-tuning induces minimal alteration to the latent space semantics. Images generated from identical noise points retain consistent content, with a slight improvement in reconstruction fidelity (e.g., the 9th to 11th images from the end in the final row).
D.2 robotic manipulation
These experiments were conducted under a unified framework, systematically evaluated across multiple environments. The configuration encompasses four main aspects: model architectures, training hyperparameters, environment-specific parameters, and execution modes, ensuring reproducibility and consistency throughout the study.
We employ two types of network architectures to learn the vector field: a Conditional UNet-1D and a Transformer-based diffusion model. The UNet model was designed with an input dimension matching the action space of the environment and an output dimension corresponding to the action dimensions. It incorporated global conditioning based on visual feature encodings, a hidden size of 256, 4 UNet blocks with 2 layers each, and attention mechanisms to enhance representational capacity. The Transformer model used a hidden dimension of 512, 6 transformer layers, 8 attention heads, and a dropout rate of 0.1 to prevent overfitting. Both models utilized a ResNet-18 vision encoder for extracting image features, with GroupNorm substituted for BatchNorm to improve training stability.

### v2

@@ A.3.3 Details and Proof of the Theorem
It follows from the Lipschitz continuity Assumption A.8 that $|p_{j}^{i}(x,\xi)|=|f_{j}^{i}((W_{j}x)^{i}+(W_{j}\xi)^{i})-f_{j}^{i}((W_{j}x)^{i})|\leq L_{j}^{i}|(W_{j}\xi)^{i}|$. Here, the superscript $i$ denotes the $i$-th component of the vector. When $(W_{j}x)^{i}\geq 0$, we have $p_{j}^{i}(x,\xi)=f_{j}^{i}((W_{j}x)^{i}+(W_{j}\xi)^{i})-f_{j}^{i}((W_{j}x)^{i})\geq 0$ due to the Monotonicity in Assumption A.8. Then $p_{j}^{i}(x,\xi)\leq L_{j}^{i}(W_{j}\xi)^{i}$. Multiplying both sides by a non-negative number $p_{j}^{i}(x,\xi)$, we get $p_{j}^{i}(x,\xi)^{2}\leq L_{j}^{i}p_{j}^{i}(x,\xi)(W_{j}x)^{i}$. When $(W_{j}x)^{i}\leq 0$, we have $p_{j}^{i}(x,\xi)=f_{j}^{i}((W_{j}x)^{i}+(W_{j}\xi)^{i})-f_{j}^{i}((W_{j}x)^{i})\leq 0$ due to the same Monotonicity Assumption, leading to $p_{j}^{i}(x,\xi)\leq-L_{j}^{i}(W_{j}\xi)^{i}$. Multiplying both sides by a non-negative number $p_{j}^{i}(x,\xi)$, we get the same result $p_{j}^{i}(x,\xi)^{2}\leq L_{j}^{i}p_{j}^{i}(x,\xi)(W_{j}x)^{i}$.
Summing over $i$ on both sides gives $\sum_{i}p_{j}^{i}(x,\xi)^{2}\leq\sum_{i}L_{j}^{i}p_{j}^{i}(x,\xi)(W_{j}x)^{i}$. Or equivalently, in a compact from, $p_{j}(x,\xi)^{\top}p_{j}(x,\xi)\leq(W_{j}x)^{\top}L_{j}p_{j}(x,\xi)=x^{\top}W_{j}^{\top}L_{j}p_{j}(x,\xi)$. This completes the proof of the first part of the lemma. By noting the relation of $f(W_{j}x)=f(W_{j}x-0)=p_{j}(0,\xi)$, the second part of the lemma holds naturally. ∎
We now begin the proof of the theorem.
Proof of Theorem A.10.
@@
Then, under Lemma A.11, it can be deduced that
@@
Therefore, with the conditions (A.11), we can substantiate that the error dynamics of system (A.8) asymptotically approaches zero, meaning that the solution is contracting. Moreover, by Khalil and Grizzle (2002, Theorem 4.9). the stability area of $\xi$ contains $\{\xi|V(\xi)\leq\max_{\zeta\in\mathbb{R}^{d}}\tilde{V}(\zeta)\}$. Since $y(t)=x(t)+\xi(t)$, we can determine that $\forall y_{0}\in\{x_{0}+\xi|V(\xi)\leq\max_{\zeta\in\mathbb{R}^{d}}\tilde{V}(\zeta)\}$, $y(t)\to x(t)$. This completes the proof. ∎
Remark A.12.
Theorems A.9 and A.10 establish the ISS and contraction property for systems of the form (A.8) via tractable linear inequalities. For the contraction analysis of these more general forms, we can refer to Lohmiller and Slotine (1998) and Li et al. (2025), albeit at the expense of increased condition complexity.
A.4 Corollary 4.3 and the Direct Parameterization Scheme
We first provide the proof of Corollary 4.3.
Proof of Corollary 4.3.
The proof is straightforward. After a simple transformation, $\delta(x)=A_{0}x(t)+\sum_{j=1}^{M}A_{j}f_{j}(D_{j}x(t))=A_{0}x(t)+\sum_{j=1}^{M}A_{j}\hat{f}_{j}(x(t))$, where $\hat{f}_{j}^{i}=D_{j}^{(ii)}f_{j}^{i}(x^{i}(t))$. Given that all entries of the diagonal matrix $D_{j}$ are positive, it is evident that $\hat{f}$ also satisfies Assumptions A.7-A.8. The proof follows directly from the results of Mei et al. (2022). ∎
Direct Parameterization.
Next, we detail how to leverage Corollary 4.3 to construct a direct parameterization scheme. Note that we can construct a Hurwitz matrix via the Schur decomposition. Assuming all eigenvalues of $A_{j}$ are real, since we don’t seem to have the need to use complex eigenvalues to characterize oscillations in generative models. To be more specific, we let $A_{j}=Q_{j}T_{j}Q_{j}^{T}$, where $T$ is an upper triangular matrix with negative diagonal elements, and $Q_{j}$ is an orthogonal matrix. We further use the Cayley transform to parameterize $Q_{j}$ as $Q_{j}=(I+K_{j})(I-K_{j})^{-1}$, where $I$ is the identity matrix and $K_{j}$ is an arbitrary skew-symmetric matrix ($K_{j}^{T}=-K_{j}$), which has d(d1)/2d(d-1)/2 degrees of freedom. The computational complexity of matrix inversion in the Cayley transform can be addressed by using convolutions in the Fourier domain (Trockman and Kolter, 2021). However, here we adopt a simpler and more straightforward strategy that relies on a low-rank assumption on $K_{j}$. Let $K_{j}=ef^{T}-fe^{T}$, where $e,f\in\mathbb{R}^{d}$. Then using the Sherman-Morrison-Woodbury (SMW) formula, we have $(I-K)^{-1}=(I-ECF^{T})^{-1}$, where $E=[e,f]\in\mathbb{R}^{d\times 2},F=[f,e]\in\mathbb{R}^{d\times 2},C=[0,-1;1,0]\in\mathbb{R}^{2\times 2}$. Note that here we only require the inversion of a two-dimensional matrix. The previously involved positive quantities can be obtained using $\exp(\cdot)$ or $\operatorname{softplus}(\cdot)$ functions. Thus, we have implemented a direct parameterization scheme that inherently satisfies conditions in Theorem 4.3 during training.
A.5 Algorithm Workflow
Algorithm workflows of FM, MLE finetuning and residual finetuning with generalized artificial viscosity are presented in Algorithms A.1-A.3.
Algorithm A.1 Flow Matching Training
0: observation $\bm{O}$, samples $\bm{X}_{1}$, source random waypoints $p_{0}$
0: flow $\mathbf{v}_{\bm{\theta}}$
1:while not converged do
2:  $\bm{x}_{0}\sim p_{0}$; $\bm{x}_{1},\bm{o}\sim\bm{X}_{1},\bm{O}$, Sample a batch
3:  (Optional) reassign the pairing between $\bm{x}_{0}$ and $\bm{x}_{1}$
4:  $t\sim\mathcal{U}[0,1]$, sample time steps
5:  $\bm{x}_{t}=t\bm{x}_{1}+(1-t)\bm{x}_{0}$, linear interpolation
6:  $\nabla_{\bm{\theta}}\left\|v_{\bm{\theta}}(\bm{x}_{t},t|\bm{o})-\underbrace{\dot{\bm{x}}_{t}}_{x_{1}-x_{0}}\right\|$, gradient step
7:end while
8: Stopping criteria: training epochs reached
Algorithm A.2 MLE Finetuning
0: observation $\bm{O}$, samples $\bm{X}_{1}$, source random waypoints $p_{0}$
0: flow $\mathbf{v}_{\bm{\theta}}$
1:while not converged do
2:  $\bm{x}_{0}\sim p_{0}$; $\bm{x}_{1},\bm{o}\sim\bm{X}_{1},\bm{O}$, Sample a batch
3:  (Optional) reassign the pairing between $\bm{x}_{0}$ and $\bm{x}_{1}$
4:  Numerical Solve $\frac{d}{dt}\phi_{t}=v_{\bm{\theta}}(\bm{\phi}_{t},t|\bm{o})$ with $\phi_{0}=x_{0}$, get $\hat{\phi}_{t=1}$.
5:  $\nabla_{\bm{\theta}}\left\|\hat{\phi}_{t=1}-x_{1}\right\|$, gradient step
6:end while
7: Stopping criteria: training epochs reached
Algorithm A.3 MLE Finetuning with GAV
0: Condtion $\bm{O}$, samples $\bm{X}_{1}$, source random waypoints $p_{0}$. Hyperparamters: a positive integer $M$, nonlinear functions $f_{1},\cdots,f_{M}$, nonnegative activation function $\sigma(\cdot)$ (e.g., $\operatorname{exp}(\cdot)$). Initial weights: $\tilde{\theta}$, $\tilde{D}_{j},e_{j},f_{j}\in\mathbb{R}^{d}$, $\tilde{T_{j}}\in\mathbb{R}^{d(d-1)/2}$,
0: flow $\mathbf{v}_{\bm{\theta}}$
[…]
6:while not converged do
7:  $\bm{x}_{0}\sim p_{0}$; $\bm{x}_{1},\bm{o}\sim\bm{X}_{1},\bm{O}$, Sample a batch
8:  (Optional) reassign the pairing between $\bm{x}_{0}$ and $\bm{x}_{1}$
9:  Numerical Solve $\frac{d}{dt}\phi_{t}=v_{t}(\phi_{t};\theta)$ with initial value $\phi_{0}=x_{0}$ for $t\in[0,1]$, get $\hat{\phi}_{t=1}(x_{0})$
10:  Numerical Solve $\frac{d}{dt}\phi_{t}=\tilde{v}_{t}(\bm{o};\tilde{\theta})+\delta(\phi_{t}(x))$ with initial value $\phi_{1}=\hat{\phi}_{t=1}(x_{0})$ for $t\in[1,2]$, get $\hat{\phi}_{t=2}(x_{0})$
11:  Compute Loss $\left\|\hat{\phi}_{t=2}(x_{0})-x_{1}\right\|$, gradient step
12:end while
13: Stopping criteria: training epochs reached
Appendix B Mathematical Analysis of Viscous Term as Negative Feedback
From PDE to ODE System via Spatial Discretization
For simplicity, we take the one-dimensional viscous Burgers equation with periodic boundary conditions as an example:
$\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}=\nu\frac{\partial^{2}u}{\partial x^{2}},\quad x\in[0,L],\quad\nu>0$
Discretizing the domain into $N$ equally spaced points with spacing $\Delta x=L/N$, we define the state vector $\mathbf{u}(t)=[u_{0}(t),u_{1}(t),\dots,u_{N-1}(t)]^{T}\in\mathbb{R}^{N}$. Using central differences for both the convective and viscous terms, we obtain the semi-discrete system:
$\frac{d\mathbf{u}}{dt}=\mathbf{f}(\mathbf{u})+\nu\mathbf{A}\mathbf{u}$
where $\mathbf{f}(\mathbf{u})$ represents the discretized nonlinear convection and $\nu\mathbf{A}\mathbf{u}$ is the discretized viscous term. The matrix $\mathbf{A}$ has the circulant tridiagonal structure:
$\mathbf{A}=\frac{1}{(\Delta x)^{2}}\begin{bmatrix}-2&1&0&\cdots&0&1\\ 1&-2&1&\cdots&0&0\\ 0&1&-2&\cdots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 1&0&0&\cdots&1&-2\end{bmatrix}$
Eigenvalue Analysis Revealing Negative Feedback Structure
The eigenvalues of $\mathbf{A}$ can be computed analytically due to its circulant structure:
$\lambda_{k}(\mathbf{A})=-\frac{4}{(\Delta x)^{2}}\sin^{2}\left(\frac{\pi k}{N}\right),\quad k=0,1,\dots,N-1$
Crucially, all eigenvalues have non-positive real parts:
  • $\lambda_{0}=0$ corresponds to the constant mode (due to mass conservation with periodic boundaries)
  • $\lambda_{k}<0$ for $k=1,2,\dots,N-1$, with $\lambda_{N/2}=-4/(\Delta x)^{2}$ being the most negative for even NN
This eigenvalue structure has profound implications. In Fourier space, the system decouples into modal equations:
$\frac{d\hat{u}_{k}}{dt}=\hat{f}_{k}(\hat{\mathbf{u}})-\underbrace{|\nu\lambda_{k}|}_{\text{feedback gain}}\hat{u}_{k},\quad k=1,\dots,N-1$
The viscous term appears explicitly as linear negative feedback with gain $g_{k}=|\nu\lambda_{k}|=\frac{4\nu}{(\Delta x)^{2}}\sin^{2}(\pi k/N)$. The gain increases with wavenumber $k$, meaning higher frequency modes experience stronger damping—a characteristic feature of viscous dissipation.
Energy Dissipation as Consequence of Negative Feedback
Define the discrete kinetic energy $E(t)=\frac{\Delta x}{2}\|\mathbf{u}(t)\|^{2}$. Its time derivative is:
$\frac{dE}{dt}=\Delta x\,\mathbf{u}^{T}\frac{d\mathbf{u}}{dt}=\Delta x\left[\mathbf{u}^{T}\mathbf{f}(\mathbf{u})+\nu\mathbf{u}^{T}\mathbf{A}\mathbf{u}\right]$
The convective term conserves energy ($\mathbf{u}^{T}\mathbf{f}(\mathbf{u})=0$ for periodic boundaries), while the viscous term gives:
$\mathbf{u}^{T}\mathbf{A}\mathbf{u}=-\frac{1}{(\Delta x)^{2}}\sum_{j=0}^{N-1}(u_{j+1}-u_{j})^{2}\leq 0$
Thus,
$\frac{dE}{dt}=-\frac{\nu}{\Delta x}\sum_{j=0}^{N-1}(u_{j+1}-u_{j})^{2}\leq 0$
The viscous term is the sole energy sink, and the negativity of $\mathbf{u}^{T}\mathbf{A}\mathbf{u}$ directly follows from the negative definiteness of $\mathbf{A}$ on the space of non-constant modes.
Control-Theoretic Interpretation
From a control perspective, the discretized viscous term implements distributed state feedback:
$\text{Viscous term}=\nu\mathbf{A}\mathbf{u}=\underbrace{-\frac{2\nu}{(\Delta x)^{2}}\mathbf{I}}_{\text{local damping}}\mathbf{u}+\underbrace{\frac{\nu}{(\Delta x)^{2}}\mathbf{B}}_{\text{neighbor coupling}}\mathbf{u}$
where $\mathbf{B}$ has zeros on the diagonal and ones on the sub/super-diagonals (with periodic wrapping). The first term provides direct negative feedback proportional to each grid point’s own velocity, while the second term couples neighboring points. The overall effect drives each $u_{j}$ toward the average of its neighbors:
$(\mathbf{A}\mathbf{u})_{j}\propto\left(\frac{u_{j-1}+u_{j+1}}{2}-u_{j}\right)$
This is a distributed control law that minimizes velocity differences between adjacent points.
Stabilizing Role Against Nonlinear Instability
Even when the nonlinear convective term $\mathbf{f}(\mathbf{u})$ may induce instability (e.g., by steepening gradients toward shock formation), the viscous term’s negative feedback provides stabilization. The linearized system around any state $\mathbf{u}^{*}$ is:
$\frac{d\delta\mathbf{u}}{dt}=\mathbf{J}_{\mathbf{f}}(\mathbf{u}^{*})\delta\mathbf{u}+\nu\mathbf{A}\delta\mathbf{u}$
where $\mathbf{J}_{\mathbf{f}}$ is the Jacobian of $\mathbf{f}$. While $\mathbf{J}_{\mathbf{f}}$ may have eigenvalues with positive real parts (indicating local instability), the addition of $\nu\mathbf{A}$—with its strictly negative eigenvalues for non-constant modes—can shift the overall system eigenvalues into the left half-plane, provided $\nu$ is sufficiently large relative to the convective time scale.
Conclusion
The spatial discretization of the viscous term in the Burgers equation reveals its fundamental nature as a linear negative feedback operator. The matrix $\mathbf{A}$ is negative semi-definite, with eigenvalues corresponding to modal damping rates. This feedback structure monotonically dissipates kinetic energy, drives the velocity field toward spatial uniformity, regularizes potential singularities arising from nonlinear convection, and inherently provides numerical stability to the discrete system. Thus, viscosity in fluid systems—whether physical or numerical—acts as a distributed controller that enforces stability by penalizing spatial velocity variations, embodying the principle that negative feedback creates order from chaos.
Appendix C Experiment Details and Additional Results
C.1 Toy Example
C.1.1 Motivation & Broader Context
The core challenge addressed in this experiment—identifying and characterizing the latent structure of multi-modal distributions from samples—is a cornerstone of unsupervised learning with profound implications across disciplines. Our goal is to develop a method that not only detects distinct modes but also robustly discriminates them from minor variations, akin to learning a parsimonious deterministic model from stochastic demonstrations. This capability is critical in diverse scenarios where the observed data is a mixture generated by several latent sources or decisions:
  • Geology & Resource Exploration: In mineral prospecting, the observed data are the spatial locations of ore fragments (e.g., gold grains) found in sediment samples. The underlying generative process is a mixture: each active deposit acts as a latent center, emitting fragments with a dispersion pattern that can be modeled by a spatial distribution (e.g., a 2D Gaussian centered at the deposit). The collected sample thus forms a multi-modal spatial point pattern, where each mode’s center indicates a potential deposit location, and the spread indicates erosion and transport effects. The key inference tasks are to determine how many deposits exist and where they are located, directly mirroring the problem of estimating the number and parameters of components in a Gaussian Mixture Model (GMM).
  • Ecology & Wildlife Conservation: Data from camera traps or GPS collars provide sporadic location points of individuals from one or more animal populations. The distribution of these points in a landscape is intrinsically multi-modal. Each population has a preferred home range or core habitat (a latent center), resulting in a higher density of observations around it, forming one mode. Multiple populations, possibly of the same species with separate territories, give rise to multiple modes. The challenge is to deconvolve this mixture to estimate the number of distinct groups and the central point of each group’s range, separating true population centers from random individual movement or outlier sightings.
  • Climate Science & Atmospheric Regimes: Daily weather maps (e.g., sea-level pressure fields) represent points in a high-dimensional state space. Over time, the system does not visit all states uniformly but tends to linger around a few preferred, recurrent patterns known as weather regimes or teleconnection patterns (e.g., the North Atlantic Oscillation). Each regime acts as an attractor, making states near it more probable and forming a density cluster. The long-term climate record thus represents a mixture of these regimes. Identifying these regimes involves clustering in state space to find the number of recurrent patterns (modes) and their characteristic spatial structures (centroids). This is crucial for understanding low-frequency climate variability and improving sub-seasonal forecasting.
  • Imitation Learning & Robotic Decision-Making: Consider an autonomous vehicle learning obstacle avoidance from human driving data. When an obstacle appears ahead, different drivers may choose entirely different strategies — such as turning left (Mode A), turning right (Mode B), or braking to a stop (Mode C) — resulting in a multimodal distribution in the action space. A simple behavioral cloning model trained to minimize average error would converge toward the average of these modes, likely producing an unsafe, indecisive response, such as hesitant steering or delayed braking. Additionally, within each strategic mode — for example, all left-turn trajectories — there exists natural variation in execution, such as slight differences in steering angle or timing. A robust method must therefore distinguish between meaningful strategic modes and treat execution-level variations as noise to be filtered out rather than as separate strategies, capturing the essential challenge of extracting clear multimodal policy structure from noisy demonstration data.
In all these cases, the number of contributing sources or high-level decisions is unknown a priori. A successful method must perform dual inference: estimating this number and robustly identifying the parameters of each major component while being robust to within-mode noise. Our experimental framework is designed to rigorously evaluate this critical capacity for recovering true latent structure from mixture samples.
@@ C.1.1 Motivation & Broader Context
C.1.2 More Discussion on Reconstruction Error Optimization in Precision-Demanding Tasks
@@
C.1.3 Experimental Details
Target Distribution Configuration.
The target distribution is a 2D mixture Gaussian distribution with five centers at coordinates [0.8,0.5][-0.8,-0.5], [0.1,0.3][-0.1,-0.3], [0.6,0.4][0.6,-0.4], [0.4,0.8][-0.4,0.8], and $[0.5,0.7]$. The mixture weights are non-uniform: $[0.3,0.2,0.2,0.1,0.2]$. Each component has a standard deviation $\sigma=0.005$ with diagonal covariance matrix $\sigma^{2}I=0.0001\times I$.
Model Architecture.
We employ a FM model based on Optimal Transport Conditional Flow Matching (OT-CFM). The network is a 3-layer fully connected neural network with 128 hidden dimensions and ReLU activation functions. The model takes 2D spatial coordinates concatenated with time $t$ as input and outputs a 2D velocity field vector.
Pretraining Hyperparameters.
The model is pretrained for 2000 epochs with a batch size of 512. We use the Adam optimizer with a learning rate of 0.001 and cosine annealing learning rate scheduler. The loss function is the flow matching loss with OT pairing mode and $\sigma=0.0$ for deterministic flow matching.
MLE Fine-tuning Hyperparameters.
For MLE fine-tuning, we train for 1000 epochs with batch size 512. The learning rate is 0.001 using Adam optimizer with weight decay $1\times 10^{-7}$ and cosine annealing scheduler. The loss function is MSE loss with 10 sampling steps and sorted pairs pairing mode. Fine-tuning starts from the pretrained checkpoint at epoch 1000.
C.1.4 Additional Results
Figure C.1 illustrates the error distributions of the Euler discretization method under different step sizes (1, 2, 4, 8, 16). MLE Fintuning consistently achieved better inference results across different scales. Figure C.2 draws the inference trajectories of the pre-trained and fine-tuned models starting from some randomly selected initial points. It can be observed that after fine-tuning, the trajectory endpoints generally move closer to the center. Additionally, fine-tuning does not increase the curvature of the trajectories; in fact, some originally curved trajectories even become straighter after fine-tuning.
Figure C.1: Error visualization across the latent space: a comparison of pre-trained and fine-tuned models at different numerical precisions (Extended Version with Additional Details)
Figure C.2: trajectories
C.2 Robotic Manipulation
C.2.1 Experimental Details
We employ two types of network architectures to learn the vector field: a Conditional UNet-1D and a Transformer-based model. The UNet model was designed with an input dimension matching the action space of the environment and an output dimension corresponding to the action dimensions. It incorporated global conditioning based on visual feature encodings, a hidden size of 256, 4 UNet blocks with 2 layers each, and attention mechanisms to enhance representational capacity. The Transformer model used a hidden dimension of 512, 6 transformer layers, 8 attention heads, and a dropout rate of 0.1 to prevent overfitting. Both models utilized a ResNet-18 vision encoder for extracting image features, with GroupNorm substituted for BatchNorm to improve training stability. In the residual fine-tuning part, the normalization terms use the same architecture as in the pre-training phase, but with fewer hidden layers and parameters, and with the state input set to zero for decoupling. In the ablation study, the GAV term is replaced with a single-hidden-layer MLP comprising 128 hidden nodes. For sampling, DDPM uses 100 steps, while all other methods (DPM, DDIM, FM, etc.) employ 16 steps. Additional details can be found in the source code.